# Constrained Best Arm Identification with Tests for Feasibility

### v1

@@ INTRODUCTION
Best arm identification (BAI) contains $K$ arms that allow a decision maker to repeatedly pull one of the arms and observes an i.i.d. sample drawn from the distribution associated with that arm (bechhofer1958sequential; paulson1964sequential). In the fixed confidence setting of BAI, given a target failure rate $\delta>0$, the decision-maker aims to identify the arm with the highest expected performance, with probability at least $1-\delta$, while keeping the number of pulls to a minimum. This problem has been widely applied in areas such as drug discovery, crowdsourcing, distributed systems, and A/B testing (jun2016top; koenig1985procedure; schmidt2006integrating; van2017automatic; zhou2014optimal).
In many real-world applications, we additionally wish to find the optimal arm that satisfies feasibility constraints, which cannot be modeled analytically and need to be evaluated via costly experimentation. Moreover, these feasibility experiments can usually be tested separately from the performance evaluation. As a motivating example, in drug discovery, we wish to find the most potent drug, but also need to guarantee that the drug is soluble in the bloodstream, and the risk of adverse side effects is small. Scientists have developed different experiments to measure the drug’s potency (performance), solubility (feasibility), and toxicity (feasibility), which can be separately carried out independent of each other (thall1998strategy). Similarly, in database tuning applications (van2017automatic), we wish to minimize the end-to-end latency, while guaranteeing the risk of system-wide failures is small. The latency and robustness can also be evaluated via separate tests (kanellis2020too).
Majority of the prior work on BAI does not study feasibility constraints. The only work on feasible BAI (katz2018feasible; katz2019top) assumes that the performance and feasibility tests are all conducted simultaneously, which does not hold in many real-world settings, such as the examples highlighted above Often, we can conduct separate experiments and only observe the result of the experiment conducted. Thus, applying methods in prior works to our setting by naively pulling all arms simultaneously can be unnecessarily expensive as it samples all performance and feasibility distributions when testing one arm and does not focus on the most important experiment to determine an arm’s optimality and/or feasibility. Moreover, from a theoretical perspective, simultaneous pulls cannot capture the actual complexity of the number of individual tests.
@@
The bandit framework is a popular paradigm to study exploration-exploitation tradeoffs that occur in sequential decision-making under uncertainty (lai85bandits; thompson33sampling). Here, a decision-maker adaptively samples one of $K$ arms from a bandit model so as to achieve a given objective. There is extensive prior work on developing algorithms for BAI, whose goal is to identify the arm with the highest mean in the fixed confidence or fixed budget setting (bechhofer1958sequential; bubeck2009pure; even2002pac; jamieson2014lil; kalyanakrishnan2012pac; karnin2013almost; paulson1964sequential). Several works have also developed hardness results for BAI (kaufmann2016complexity; mannor2004sample). Our algorithm and theoretical results build on this rich line of work. In particular, our algorithm’s sampling strategy is inspired by the LUCB algorithm of (kalyanakrishnan2012pac). However, none of these work studies constrained BAI, where the feasibility constraints need to be tested separately.
@@
Our model for testing feasibility is closely related to thresholding bandits, i.e., identify all arms whose mean is larger than a given threshold (katz2018feasible; locatelli2016optimal; mason2020findall). In our work, however, it is sufficient to identify just one constraint that is larger than the threshold to determine if it is infeasible. Another similar line of work is on finding if there is any arm below a threshold (degenne2019pure; kano2019good; katz2020true; kaufmann2018sequential). While we use some ideas from these works when testing the feasibility of an arm, their algorithms and analysis cannot be directly applied to our setting, since we need to determine if the most efficient way to eliminate an arm is via its sub-optimality or infeasibility.
@@
Perhaps the closest work to ours is (katz2019top), who study identifying the best $m$ arms that satisfy the given feasibility constraints. They assume each arm $i$ is associated with a $D$-dimensional distribution with mean $\bm{\mu}_{i}$ and aim to find the top $m$ arms that maximize $\mathbf{r}^{\top}\bm{\mu}_{i}$ subject to the constraint $\bm{\mu}_{i}\in P$. Here $\mathbf{r}$ is a given direction for the reward and $P$ is a subset of $\mathbb{R}^{D}$ denoting the feasible set. However, this is different from our setting, since we allow the objective and feasibility constraints to be tested separately. A naive application of their algorithm will require testing an arm’s performance and all feasibility constraints each time when we wish to test an arm, which is sample inefficient in our setting. chen2017nearly studies identifying the best set in a family of feasible subsets. Our work is different from theirs since they did not consider testing for feasibility separately and they assume the feasible subsets have various combinatorial structures.
@@
Our work is also related to work in the Bayesian optimization literature that studies zeroth order constrained optimization and multi-objective optimization, where multiple performance objectives and constraints can be evaluated separately, similar to our setting (gardner2014bayesian; ungredda2021bayesian; eriksson2021scalable; kirschner2022tuning; hernandez2016; berkenkamp2021bayesian; kirschner2019adaptive; paria2020flexible; hernandez2016predictive). These problem settings are distinctly different from ours as we consider the KK–armed version of the problem.
@@
Inspired from kano2019good, we design the subroutine SAMPLE-FEASIBILITY, which recommends one constraint to test when we wish to test the feasibility of an arm. Each time we invoke it with an arm $i\in[K]$, it chooses the constraint in $H_{i}$ that maximizes $\widetilde{\mu}_{i,\ell}(t)$, as in (6). Intuitively, via this choice, we are sampling the constraint that is most likely to be larger than the threshold $\frac{1}{2}$. After sampling, if the lower confidence bound for the sampled constraint is larger than $\frac{1}{2}$, we deem it to be infeasible so we add $i$ to $I$; then $i$ will be eliminated from $S$ shortly. otherwise, if its upper confidence bound is smaller than $\frac{1}{2}$, we have verified that the sampled constraint is not in violation and remove it from $H_{i}$. In the event that $H_{i}$ is empty, then none of the constraints are in violation, and hence we add $i$ to $F$ as we have verified that $i$ is feasible.
@@
Gap between the lower and upper bound when $\delta\nrightarrow 0$: It is easy to verify that the sample complexity of Algorithm 1 matches the lower bound in Theorem 1 when $N=1$ for any $\delta\in(0,1)$. However, when $N>1$, Algorithm 1 can still match the lower bound for any $\delta$ on all arms except for arms in $\mathcal{I}$. This mismatch is due to that Theorem 1 can only be achieved by an oracle who already knows the structure and the means of all arms and samples just to verify that $i^{\star}$ is the optimal arm. Theorem 1 suggests that we should only sample the constraint that has the highest mean for arms in $\mathcal{I}$. However, since we assume the decision-maker does not have all the additional information, it is expected to invest in samples on other constraints (not the highest one) for exploration. This also explains why in Thoerem 2, we have the lower order terms ($G_{1},\cdots,G_{5}$), as they account for the cost for exploration. Previous works (simchowitz2017simulator; karnin2016verification) have discussed the difference between verification and exploration but their tools cannot be directly applied to our problem due to the definition of feasibility. kaufmann2018sequential provides a tighter lower bound when the distributions of the arms come from an exponential family, however, it still has a large gap between their upper bound. It remains an open question to lower bound the sample complexity any algorithm need to spend on the cost for exploration.
@@
    Feasibility-First (F-first): for each arm $i$, we first use the subroutine SAMPLE-FEASIBILITY, which samples one constraint at each round, to determine whether it is feasible. Then, we apply LUCB algorithm (kalyanakrishnan2012pac) to find the best arm in the feasible set.
@@
    Performance-First (P-first): We first run LUCB algorithm (kalyanakrishnan2012pac) on all arms’ performance and then test the best arm’s feasibility using the subroutine SAMPLE-FEASIBILITY. If the best arm is feasible, it stops, otherwise it eliminates this arm and repeats the same procedure on the remaining arms.
@@
    TF-LUCB-C (katz2019top): This procedure was developed for feasible BAI, but assume the performance and feasibility constraints can be sampled together. We implement the most natural adaptation of it to our setting: pull all performance and feasibility distributions together using $N+1$ samples.
@@
    Naive BAI (Naive): It adapts from the Racing algorithm (even2002pac), which is optimal for BAI. On each round, it eliminates the arms that are considered infeasible or have lower performance than arms that are already considered feasible. It repeatedly tests the performance and all feasibility constraints of all survival arms. It stops when only one feasible arm is left or no arm is considered feasible.
@@
Drug Discovery: We investigate finding the most effective drug satisfying different feasibility constraints. The data comes from Tables 2 & 3 in genovese2013efficacy. Each arm corresponds to a dosage level (in mg): [25, 75, 150, 300, Placebo] and has $3$ attributes: probability of being effective $\mu_{i,0}$, probability of any adverse event $\mu_{i,1}$, and probability of causing an infection or infestation $\mu_{i,2}$. The threshold for constraint 1 is 0.5 and for constraint 2 is 0.25. We manually increase 0.25 to all rewards received from constraint 2 to make the thresholds equal. The means for different arms are $(.34,.519,.259)^{T}$, $(.469,.612,.184)^{T}$, $(.465,.465,.209)^{T}$, $(.537,.61,.293)^{T}$, and $(.36,.58,.16)^{T}$. We use Gaussian distribution with variance $1$. Fig. 3 shows that our algorithm outperforms all other methods.
In all the experiments, all methods can find $i^{\star}$ correctly. Moreover, Algorithm 1, P-first, and F-first, which all use SAMPLE-FEASIBILITY, generally perform better than the other two methods which sample all constraints simultaneously, demonstrating that we should test each constraint individually. Since different confidence bounds will affect the performance of the algorithm (jamieson2014lil), for a fair comparison, we use (5) on all methods. In the appendix, To see the influence of the confidence bounds on the methods, we compare Algorithm 1 and TF-LUCB-C with two different confidence bounds to see the influence of different confidence bounds. It shows that our algorithm outperforms TF-LUCB-C in both bounds and even when only TF-LUCB-C uses a tighter confidence bound.
@@ CONCLUSION
The main challenge in the proof of this lower bound (with an asymptotically matching upper bound) is in defining the complexity terms in Section LOWER BOUND. Our proof essentially considers different cases, and then applies the change-of-measure lemma from kaufmann2016complexity. Although this lemma was presented for the case where each arm only contains one distribution, it easily generalizes to our setting where we have $K(N+1)$ distributions, as we assume no dependence between the distributions. For completeness, we have stated this result in Lemma 19 in Appendix Some useful results.
@@
Lower bounding $d(\mathbb{P}_{\nu}(\mathcal{A}),\mathbb{P}_{\nu^{\prime}}(\mathcal{A}))\geq\log\frac{1}{2.4\delta}$ as in kaufmann2016complexity and letting $\alpha\to 0$, we conclude that
@@
This proof is organized as follows. First, in Lemma 4 we will define a “clean” event, and show that this event holds with probability at least $1-\delta$. Using this result, in Lemma 3, we will show that the algorithm always identifies the correct $i^{\star}$ upon termination under this clean event. This technique of proving an algorithm is $\delta$-correct is widely adopted in the bandit literature (wang2022best; kano2019good).
@@
The main challenge in obtaining an upper bound that matches the lower bound requires a careful decomposition of the stopping time $\tau$, i.e. of the number of epochs of Algorithm 1 upon termination. We decompose $\tau$ into two parts based on whether the sampled arms $a_{t},b_{t}$ belong to $\mathcal{I}$ or not. Furthermore, in each part, if an arm $i$ is chosen as in line 10 in Algorithm 1, we will treat $a_{t}=i$ and $b_{t}$ as undefined. This decomposition helps us to consider the two different types of elimination (feasibility and worse performance) separately. To eliminate an arm by performance, we leverage the insight that if two arms are sampled at an epoch, the confidence intervals of the two arms must overlap. We then upper bound the time for an arm to be eliminated via performance by the sample complexity to differentiate this arm’s performance from other arms. To eliminate an arm by feasibility, We adapt the analysis from kano2019good to upper bound the expected sample complexity of determining an arm’s feasibility. The insight here is that if a constraint of an arm is not determined at an epoch, this constraint’s confidence interval must contain the threshold 1/2.
@@
For part $B_{7}$, this indicates that all feasible arms are not wrongly identified as infeasible. This is the standard BAI problem. Note that even if some arms $i$ in $\{i\in\mathcal{F}^{\mathsf{c}}:\theta_{i}>\phi_{i}\}$ are identified as infeasible in this scenario, they will only be eliminated faster which means $B_{7}$ is a bit looser for these arms. Thus we apply the result of the sample complexity of the LUCB Algorithm (kalyanakrishnan2012pac) to upper bound $B_{7}$. Theorem 6 from kalyanakrishnan2012pac indicates the expected sample complexity for the standard BAI problem using the LUCB algorithm, thus we have
@@
We will first discuss the case when the arm is not in $\mathcal{F}$ in Lemma 6, and then discuss the case when the arm is in $\mathcal{F}$ in Lemma 7. For this proof, we borrow some analysis techniques from kano2019good who study a version of the thresholding bandits problem and adapt them to our setting.
@@
Recall that $N_{i,\ell}(r)$ denotes the number of times $(i,\ell)$ has been pulled up to round $r$. In particular, Lemma 1 in kaufmann2016complexity holds as:
@@
We follow the remark in kaufmann2016complexity that

### v2

@@ INTRODUCTION
Best arm identification (BAI) contains $K$ arms that allow a decision maker to repeatedly pull one of the arms and observes an i.i.d. sample drawn from the distribution associated with that arm (Bechhofer 1958; Paulson 1964). In the fixed confidence setting of BAI, given a target failure rate $\delta>0$, the decision-maker aims to identify the arm with the highest expected performance, with probability at least $1-\delta$, while keeping the number of pulls to a minimum. This problem has been widely applied in areas such as drug discovery, crowdsourcing, distributed systems, and A/B testing (Jun et al. 2016; Koenig and Law 1985; Schmidt et al. 2006; Van Aken et al. 2017; Zhou et al. 2014).
In many real-world applications, we additionally wish to find the optimal arm that satisfies feasibility constraints, which cannot be modeled analytically and need to be evaluated via costly experimentation. Moreover, these feasibility experiments can usually be tested separately from the performance evaluation. As a motivating example, in drug discovery, we wish to find the most potent drug, but also need to guarantee that the drug is soluble in the bloodstream, and the risk of adverse side effects is small. Scientists have developed different experiments to measure the drug’s potency (performance), solubility (feasibility), and toxicity (feasibility), which can be separately carried out independent of each other (Thall and Russell 1998). Similarly, in database tuning applications (Van Aken et al. 2017), we wish to minimize the end-to-end latency, while guaranteeing the risk of system-wide failures is small. The latency and robustness can also be evaluated via separate tests (Kanellis et al. 2020).
Majority of the prior work on BAI does not study feasibility constraints. The only work on feasible BAI (Katz-Samuels and Scott 2018, 2019) assumes that the performance and feasibility tests are all conducted simultaneously, which does not hold in many real-world settings, such as the examples highlighted above Often, we can conduct separate experiments and only observe the result of the experiment conducted. Thus, applying methods in prior works to our setting by naively pulling all arms simultaneously can be unnecessarily expensive as it samples all performance and feasibility distributions when testing one arm and does not focus on the most important experiment to determine an arm’s optimality and/or feasibility. Moreover, from a theoretical perspective, simultaneous pulls cannot capture the actual complexity of the number of individual tests.
@@
The bandit framework is a popular paradigm to study exploration-exploitation tradeoffs that occur in sequential decision-making under uncertainty (Lai and Robbins 1985; Thompson 1933). Here, a decision-maker adaptively samples one of $K$ arms from a bandit model so as to achieve a given objective. There is extensive prior work on developing algorithms for BAI, whose goal is to identify the arm with the highest mean in the fixed confidence or fixed budget setting (Bechhofer 1958; Bubeck et al. 2009; Even-Dar et al. 2002; Jamieson et al. 2014; Kalyanakrishnan et al. 2012; Karnin et al. 2013; Paulson 1964). Several works have also developed hardness results for BAI (Kaufmann et al. 2016; Mannor and Tsitsiklis 2004). Our algorithm and theoretical results build on this rich line of work. In particular, our algorithm’s sampling strategy is inspired by the LUCB algorithm of (Kalyanakrishnan et al. 2012). However, none of these work studies constrained BAI, where the feasibility constraints need to be tested separately.
@@
Our model for testing feasibility is closely related to thresholding bandits, i.e., identify all arms whose mean is larger than a given threshold (Katz-Samuels and Scott 2018; Locatelli et al. 2016; Mason et al. 2020). In our work, however, it is sufficient to identify just one constraint that is larger than the threshold to determine if it is infeasible. Another similar line of work is on finding if there is any arm below a threshold (Degenne and Koolen 2019; Kano et al. 2019; Katz-Samuels and Jamieson 2020; Kaufmann et al. 2018). (Hou et al. 2023) consider a variance-constrained BAI problem, where feasibility is defined via a variance threshold and is not tested separately from performance, as both are derived from the same reward distribution. While we use some ideas from these works when testing the feasibility of an arm, their algorithms and analysis cannot be directly applied to our setting, since we need to determine if the most efficient way to eliminate an arm is via its sub-optimality or infeasibility.
@@
Perhaps the closest work to ours is (Katz-Samuels and Scott 2019), who study identifying the best $m$ arms that satisfy the given feasibility constraints. They assume each arm $i$ is associated with a $D$-dimensional distribution with mean $\bm{\mu}_{i}$ and aim to find the top $m$ arms that maximize $\mathbf{r}^{\top}\bm{\mu}_{i}$ subject to the constraint $\bm{\mu}_{i}\in P$. Here $\mathbf{r}$ is a given direction for the reward and $P$ is a subset of $\mathbb{R}^{D}$ denoting the feasible set. However, this is different from our setting, since we allow the objective and feasibility constraints to be tested separately. A naive application of their algorithm will require testing an arm’s performance and all feasibility constraints each time when we wish to test an arm, which is sample inefficient in our setting. Chen et al. (2017) studies identifying the best set in a family of feasible subsets. Our work is different from theirs since they did not consider testing for feasibility separately and they assume the feasible subsets have various combinatorial structures.
@@
Our work is also related to work in the Bayesian optimization literature that studies zeroth order constrained optimization and multi-objective optimization, where multiple performance objectives and constraints can be evaluated separately, similar to our setting (Gardner et al. 2014; Ungredda and Branke 2021; Eriksson and Poloczek 2021; Kirschner et al. 2022; Hernández-Lobato et al. 2016b; Berkenkamp et al. 2021; Kirschner et al. 2019; Paria et al. 2020; Hernández-Lobato et al. 2016a). These problem settings are distinctly different from ours as we consider the KK–armed version of the problem.
@@
Inspired from Kano et al. (2019), we design the subroutine SAMPLE-FEASIBILITY, which recommends one constraint to test when we wish to test the feasibility of an arm. Each time we invoke it with an arm $i\in[K]$, it chooses the constraint in $H_{i}$ that maximizes $\widetilde{\mu}_{i,\ell}(t)$, as in (6). Intuitively, via this choice, we are sampling the constraint that is most likely to be larger than the threshold $\frac{1}{2}$. After sampling, if the lower confidence bound for the sampled constraint is larger than $\frac{1}{2}$, we deem it to be infeasible so we add $i$ to $I$; then $i$ will be eliminated from $S$ shortly. otherwise, if its upper confidence bound is smaller than $\frac{1}{2}$, we have verified that the sampled constraint is not in violation and remove it from $H_{i}$. In the event that $H_{i}$ is empty, then none of the constraints are in violation, and hence we add $i$ to $F$ as we have verified that $i$ is feasible.
@@
Gap between the lower and upper bound when $\delta\nrightarrow 0$: It is easy to verify that the sample complexity of Algorithm 1 matches the lower bound in Theorem 1 when $N=1$ for any $\delta\in(0,1)$. However, when $N>1$, Algorithm 1 can still match the lower bound for any $\delta$ on all arms except for arms in $\mathcal{I}$. This mismatch is due to that Theorem 1 can only be achieved by an oracle who already knows the structure and the means of all arms and samples just to verify that $i^{\star}$ is the optimal arm. Theorem 1 suggests that we should only sample the constraint that has the highest mean for arms in $\mathcal{I}$. However, since we assume the decision-maker does not have all the additional information, it is expected to invest in samples on other constraints (not the highest one) for exploration. This also explains why in Thoerem 2, we have the lower order terms ($G_{1},\cdots,G_{5}$), as they account for the cost for exploration. Previous works (Simchowitz et al. 2017; Karnin 2016) have discussed the difference between verification and exploration but their tools cannot be directly applied to our problem due to the definition of feasibility. Kaufmann et al. (2018) provides a tighter lower bound when the distributions of the arms come from an exponential family, however, it still has a large gap between their upper bound. It remains an open question to lower bound the sample complexity any algorithm need to spend on the cost for exploration.
@@
    Feasibility-First (F-first): for each arm $i$, we first use the subroutine SAMPLE-FEASIBILITY, which samples one constraint at each round, to determine whether it is feasible. Then, we apply LUCB algorithm (Kalyanakrishnan et al. 2012) to find the best arm in the feasible set.
@@
    Performance-First (P-first): We first run LUCB algorithm (Kalyanakrishnan et al. 2012) on all arms’ performance and then test the best arm’s feasibility using the subroutine SAMPLE-FEASIBILITY. If the best arm is feasible, it stops, otherwise it eliminates this arm and repeats the same procedure on the remaining arms.
@@
    TF-LUCB-C (Katz-Samuels and Scott 2019): This procedure was developed for feasible BAI, but assume the performance and feasibility constraints can be sampled together. We implement the most natural adaptation of it to our setting: pull all performance and feasibility distributions together using $N+1$ samples.
@@
    Naive BAI (Naive): It adapts from the Racing algorithm (Even-Dar et al. 2002), which is optimal for BAI. On each round, it eliminates the arms that are considered infeasible or have lower performance than arms that are already considered feasible. It repeatedly tests the performance and all feasibility constraints of all survival arms. It stops when only one feasible arm is left or no arm is considered feasible.
@@
Drug Discovery: We investigate finding the most effective drug satisfying different feasibility constraints. The data comes from Tables 2 & 3 in Genovese et al. (2013). Each arm corresponds to a dosage level (in mg): [25, 75, 150, 300, Placebo] and has $3$ attributes: probability of being effective $\mu_{i,0}$, probability of any adverse event $\mu_{i,1}$, and probability of causing an infection or infestation $\mu_{i,2}$. The threshold for constraint 1 is 0.5 and for constraint 2 is 0.25. We manually increase 0.25 to all rewards received from constraint 2 to make the thresholds equal. The means for different arms are $(.34,.519,.259)^{T}$, $(.469,.612,.184)^{T}$, $(.465,.465,.209)^{T}$, $(.537,.61,.293)^{T}$, and $(.36,.58,.16)^{T}$. We use Gaussian distribution with variance $1$. Fig. 3 shows that our algorithm outperforms all other methods.
In all the experiments, all methods can find $i^{\star}$ correctly. Moreover, Algorithm 1, P-first, and F-first, which all use SAMPLE-FEASIBILITY, generally perform better than the other two methods which sample all constraints simultaneously, demonstrating that we should test each constraint individually. Since different confidence bounds will affect the performance of the algorithm (Jamieson et al. 2014), for a fair comparison, we use (5) on all methods. In the appendix, To see the influence of the confidence bounds on the methods, we compare Algorithm 1 and TF-LUCB-C with two different confidence bounds to see the influence of different confidence bounds. It shows that our algorithm outperforms TF-LUCB-C in both bounds and even when only TF-LUCB-C uses a tighter confidence bound.
@@ CONCLUSION
The main challenge in the proof of this lower bound (with an asymptotically matching upper bound) is in defining the complexity terms in Section LOWER BOUND. Our proof essentially considers different cases, and then applies the change-of-measure lemma from Kaufmann et al. (2016). Although this lemma was presented for the case where each arm only contains one distribution, it easily generalizes to our setting where we have $K(N+1)$ distributions, as we assume no dependence between the distributions. For completeness, we have stated this result in Lemma 19 in Appendix Some useful results.
@@
Lower bounding $d(\mathbb{P}_{\nu}(\mathcal{A}),\mathbb{P}_{\nu^{\prime}}(\mathcal{A}))\geq\log\frac{1}{2.4\delta}$ as in Kaufmann et al. (2016) and letting $\alpha\to 0$, we conclude that
@@
This proof is organized as follows. First, in Lemma 4 we will define a “clean” event, and show that this event holds with probability at least $1-\delta$. Using this result, in Lemma 3, we will show that the algorithm always identifies the correct $i^{\star}$ upon termination under this clean event. This technique of proving an algorithm is $\delta$-correct is widely adopted in the bandit literature (Wang et al. 2022; Kano et al. 2019).
@@
The main challenge in obtaining an upper bound that matches the lower bound requires a careful decomposition of the stopping time $\tau$, i.e. of the number of epochs of Algorithm 1 upon termination. We decompose $\tau$ into two parts based on whether the sampled arms $a_{t},b_{t}$ belong to $\mathcal{I}$ or not. Furthermore, in each part, if an arm $i$ is chosen as in line 10 in Algorithm 1, we will treat $a_{t}=i$ and $b_{t}$ as undefined. This decomposition helps us to consider the two different types of elimination (feasibility and worse performance) separately. To eliminate an arm by performance, we leverage the insight that if two arms are sampled at an epoch, the confidence intervals of the two arms must overlap. We then upper bound the time for an arm to be eliminated via performance by the sample complexity to differentiate this arm’s performance from other arms. To eliminate an arm by feasibility, We adapt the analysis from Kano et al. (2019) to upper bound the expected sample complexity of determining an arm’s feasibility. The insight here is that if a constraint of an arm is not determined at an epoch, this constraint’s confidence interval must contain the threshold 1/2.
@@
For part $B_{7}$, this indicates that all feasible arms are not wrongly identified as infeasible. This is the standard BAI problem. Note that even if some arms $i$ in $\{i\in\mathcal{F}^{\mathsf{c}}:\theta_{i}>\phi_{i}\}$ are identified as infeasible in this scenario, they will only be eliminated faster which means $B_{7}$ is a bit looser for these arms. Thus we apply the result of the sample complexity of the LUCB Algorithm (Kalyanakrishnan et al. 2012) to upper bound $B_{7}$. Theorem 6 from Kalyanakrishnan et al. (2012) indicates the expected sample complexity for the standard BAI problem using the LUCB algorithm, thus we have
@@
We will first discuss the case when the arm is not in $\mathcal{F}$ in Lemma 6, and then discuss the case when the arm is in $\mathcal{F}$ in Lemma 7. For this proof, we borrow some analysis techniques from Kano et al. (2019) who study a version of the thresholding bandits problem and adapt them to our setting.
@@
Recall that $N_{i,\ell}(r)$ denotes the number of times $(i,\ell)$ has been pulled up to round $r$. In particular, Lemma 1 in Kaufmann et al. (2016) holds as:
@@
We follow the remark in Kaufmann et al. (2016) that